# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ I Introduction
Quantum machine learning represents a highly promising realm in contemporary physics and computer science research, with far-reaching implications spanning quantum chemistry Peruzzo et al. (2014), artificial intelligence Liu et al. (2023), and even high-energy physics Andreassen et al. (2019). Nevertheless, it remains in its nascent stages of development. This is evident from the absence of a precise definition for quantum machine learning. Some describe it as the convergence of quantum computing and machine learning, wherein machine learning algorithms are executed on quantum devices. In simpler terms, it can be thought of as the quantum counterpart to classical machine learning.
@@
In recent times, artificial intelligence, exemplified by technologies like ChatGPT, has become an integral part of everyday life. It’s entirely plausible that, in the future, we will harness artificial intellegence in an even wider array of applications, including medical diagnostics, education, and aiding scientific research. Much of artificial intellegence’s success hinges on machine learning, a departure from traditional computer programs that entail crafting explicit instructions to directly solve problems Roberts et al. (2022). Machine learning models, on the other hand, are trained on real-world data stored in a dataset (often denoted as $\mathcal{D}$), acquiring the ability to tackle problems autonomously. This represents a departure, in a sense, from the conventional von Neumann model of digital computing.
@@
Researchers might initially wonder why there’s a need for quantum machine learning when its classical counterpart has demonstrated impressive performance. One of the primary rationales stems from the fact that classical machine learning relies heavily on linear algebra procedures Biamonte et al. (2017). For instance, classical machine learning can address problems like identifying the optimal hyperplane to separate two data clusters by employing matrix inversion techniques. Quantum mechanics, at its core, is inherently grounded in linear algebra, where we construct a Hilbert space ($\mathcal{H}$) of a system described by an Hermitian operator known as the Hamiltonian ($\hat{H}$) Sakurai and Napolitano (2020). Over time, it has become evident that quantum computing has the potential to dramatically enhance a computer’s problem-solving capabilities in certain specific scenarios. To illustrate, consider the task of matrix inversion again. Classical computers typically require computational efforts with a complexity of $\mathcal{O}(N\log{N})$ to accomplish this, while quantum computers can leverage algorithms like the Harrow-Hassidim-Lloyd algorithm (HHL) algorithm Harrow et al. (2009), which has a complexity of $\mathcal{O}((\log{N})^{2})$, leading to significant speedup, in certain conditions like sparsity, fault-tolerance, small condition numbers, and fast interfaces between classical and quantum processors. Although there is no known explicit realizations of HHL algorithm at the large scale at the moment, it is still an exciting direction deserves further studies, especially when quantum devices are fast developing with more and more capabilities.
@@
Due to its vague definition, quantum machine learning encompasses a broader spectrum of topics. For instance, quantum (shadow) tomography Aaronson (2018) has gained prominence, focusing on the characterization of a given quantum state by accumulating data from various measurements. This involves determining the minimum number of identical quantum state copies needed to extract sufficient information about the state’s properties. Another facet involves machine learning for quantum physics, which entails employing (perhaps classical) machine learning tools to explore various aspects of quantum physics. Additionally, some developments in both quantum algorithms and quantum hardware are often encompassed within the broader umbrella of quantum machine learning.
@@
The solution to this predicament is known as quantum error correction (QEC) code Roffe (2019). Think of QEC as a safeguard for quantum information. Typically, quantum information is lost once it’s measured, which becomes especially likely in noisy environments. However, information protected by QEC can persist if it remains undamaged within certain limits. It’s worth noting that all error correction codes have their constraints, implying that information will inevitably be lost if it’s severely damaged. Nevertheless, error correction provides a protective buffer zone against such losses.
@@
Our objective is to implement QEC codes across all quantum devices, ushering in an era of fault-tolerant quantum computing (FTQC) in the future. This trajectory parallels the history of classical computing, where, before the invention of classical error correction, scaling up and running useful algorithms on classical computers was also a formidable challenge Chang (2006). Today, we can reliably operate classical computers everyday. Given the promising advancements in quantum error correction in recent times, our optimism about the future of quantum computing remains steadfast.
@@
However, when it comes to machine learning, classical machine learning doesn’t inherently reject noise Liu et al. (2022a). The widely recognized learning algorithm, known as stochastic gradient descent, explicitly incorporates noise, and surprisingly, this noise addition actually enhances its performance. To grasp this concept, consider that noise can effectively steer us away from saddle points, offering an automatic mechanism for avoiding them. In a way, one can interpret this as machine learning’s ability to withstand and even benefit from noise. Consequently, this insight suggests that running certain machine learning algorithms on current (NISQ) quantum devices could have some significance. This prospect promises to enrich our present-day experiences significantly, especially considering the challenges associated with constructing QEC, which could take a few years to become fully integrated. Prior to entering the era of fault-tolerant quantum computing, we’ll have the opportunity to experiment with quantum devices and apply machine learning techniques, adding a vibrant dimension to our current scientific research.
@@
In addition to discussing the current applications of quantum capabilities for machine learning, we should also let our imaginations soar. The era of fault-tolerant quantum computing (FTQC) is a foreseeable future, and it’s crucial to look ahead at what challenges we can tackle with quantum machine learning on FTQC devices. One of the most renowned and valuable algorithms in this context is the Harrow-Hassidim-Lloyd (HHL) algorithm Harrow et al. (2009). Additionally, there are other algorithms that can be deployed to address a range of problems, such as principal component analysis Biamonte et al. (2017).
@@
Beside the optimistic future that quantum machine learning has, there are also a number of controversial issues with the subject. For example, some might argue that the variational quantum algorithm will not work in some circumstance. People working on quantum landscapes theory observed the famous barren plateau phenomena which leads to a kind of no-go theorem Cerezo et al. (2022a); Liu et al. (2022b). It amounts to situations that are hard to find minimum when we optimizes our objective function by gradient descent with randomized variational circuits.
@@
Debates indeed exist regarding whether quantum algorithms can consistently deliver exponential speedups. Some argue that quantum speedup is only guaranteed when dealing with quantum information. When it comes to classical information, it’s conceivable to design classical machine learning models capable of achieving comparable average prediction accuracy Tang (2019); Huang et al. (2021a). In such cases, the computational complexity difference between classical and quantum approaches may, at worst, be a modest polynomial factor. This debate underscores the importance of carefully assessing the specific problem and context when considering the potential advantages of quantum algorithms. However, the landscape shifts when the objective is to attain a low worst-case prediction error. In this scenario, it becomes feasible to achieve an exponential divergence in computational complexities between classical and quantum approaches. This underscores the potential superiority of quantum algorithms when stringent requirements for worst-case accuracy are in play. This phenomenon can be viewed as a manifestation of classical effects casting their “shadow” in the quantum realm. It should come as no surprise to individuals well-versed in quantum theory that quantum predictions tend to align with classical theory when dealing with systems possessing a high degree of freedom. This approach can also be applied to tackle a critical issue in variational quantum algorithms known as the “output problem,” often referred to as the classical shadow.
@@
Indeed, it is truly heartening and captivating to witness the evolution of quantum computing beyond celebrated algorithms like Shor’s factoring algorithm Shor (1994) and Grover search Grover (1996). Algorithms like HHL Harrow et al. (2009) offer a versatile blueprint, shedding light on how quantum computers hold the potential to deliver significant acceleration for fundamental tasks like clustering, pattern-matching, and principal component analysis. However, as Scott Aaronson aptly noted in his paper Aaronson (2015), the pursuit of these speedups will still demand significant effort and ingenuity, as nature continues to present challenges that compel us to work diligently for these advancements.
@@
We have structured our review as follows. In section II, we focus on the historical and ongoing developments in Quantum Machine Learning (QML) during the Noisy Intermediate-Scale Quantum (NISQ) era. Within this era, one of the central frameworks is the Variational Quantum Algorithm (VQA) Bharti et al. (2021); Farhi et al. (2014); Peruzzo et al. (2014), which we explore in detail in section II.1. VQA comprises four key elements: the choice of an objective function (discussed in section II.1.1), the employment of Parameterized Quantum Circuits (PQC) with adjustable parameters optimized by classical algorithms (covered in section II.1.2), measurement strategies (explored in section II.1.3), and the classical optimizer responsible for minimizing the objective function (explained in section II.1.4). Subsequently, we provide insights into the construction of the Quantum Neural Tangent Kernel (QNTK) Liu et al. (2022c) in section II.2, offering a theoretical foundation for quantum neural networks and an understanding of stochastic gradient descent dynamics from first principles. Finally, we address the issue of barren plateaus in section II.3, approaching this topic through the lens of quantum landscape theory Cerezo et al. (2022a). Additionally, we present an alternative formulation of laziness Liu et al. (2022b), providing another perspective on the same problem.
@@
In section III, our focus shifts to quantum algorithms that have the potential for exponential speedup in the Fault-Tolerant Quantum Computing (FTQC) era. We begin by introducing Quantum Phase Estimation (QPE) Chang (2006) in section III.1, where we also delve into the Quantum Principal Component Analysis (QPCA) program Lloyd (1996). Subsequently, we present a counterpoint regarding the attainability of exponential speedup for programs such as the recommendation system in section III.2. Moving on to section III.3, we introduce the pivotal Harrow-Hassidim-Lloyd (HHL) algorithm Harrow et al. (2009), which opens up possibilities for various new quantum machine learning algorithms. In section III.3.1, we illustrate a visionary perspective on the future of machine learning algorithms with the aid of Carleman linearization Liu et al. (2021) and the HHL algorithm. Finally, in section III.4, we introduce Quantum Random Access Memory (QRAM), focusing on theoretical designs Hann et al. (2019a), practical implementation efforts Hann (2021), and its necessity in certain algorithms Biamonte et al. (2017).
@@
In section IV, we delve into various topics that amalgamate quantum principles with statistical learning theory. Our primary focus is on shadow tomography Aaronson (2018) in section IV.1, where we explore the motivation behind shadow tomography and delve into the construction of the theorem. A pivotal subject inspired by shadow tomography, the classical shadow formalism Elben et al. (2022), is thoroughly discussed in section IV.2. Moving on to section IV.2.1, we examine the application of classical shadow as an efficient quantum-to-classical information converter Huang et al. (2021a). In section IV.3, we then shift our attention to the applications of Quantum Machine Learning (QML) in the study of quantum data and quantum simulators Biamonte et al. (2017).
@@
II Noisy intermediate-scale quantum (NISQ) era
The concept of the NISQ era, which stands for Noisy Intermediate-Scale Quantum, was introduced by John Preskill in 2018 Preskill (2018). It’s essential to understand that NISQ is primarily a hardware-oriented definition and doesn’t inherently imply a specific time frame. In NISQ devices, quantum circuits are implemented, where all gates adhere to a predefined graph $G$, with qubits represented by nodes in this graph. These gates typically operate on one or two qubits. Due to the noise associated with each gate operation, NISQ algorithms are naturally constrained to shallow circuit depths Barak and Marwaha (2021).
On the other hand, “near-term” algorithms are those designed for quantum devices expected to be available in the next few years, and this term doesn’t explicitly refer to the absence of Quantum Error Correction (QEC). The notion of what constitutes the “near-term” can be subjective because different researchers may have varying opinions on how many years can be considered “near-term.” Predicting experimental progress in quantum computing is always challenging, and such predictions can be influenced by human biases. Algorithms developed for near-term hardware may become unfeasible if hardware advancements do not align with the algorithm’s experimental requirements Bharti et al. (2021).
The majority of existing NISQ algorithms leverage quantum computers’ capabilities through a hybrid quantum-classical approach. These algorithms involve offloading the classically challenging aspects of certain computations to a quantum computer while performing the remaining tasks on a suitably powerful classical device. They operate by iteratively adjusting the parameters of a parametrized quantum circuit, making them known as Variational Quantum Algorithms (VQA) (Endo et al. (2021); Cerezo et al. (2022b)).
II.1 Variational quantum algorithm (VQA)
The concept of Variational Quantum Algorithms (VQA) was initially introduced to address certain quantum chemistry challenges, notably through the well-known Variational Quantum Eigensolver (VQE) McClean et al. (2016); Peruzzo et al. (2014); Wecker et al. (2015). Additionally, VQA includes the Quantum Approximate Optimization Algorithm (QAOA) Farhi et al. (2014), which is designed for solving combinatorial optimization problems.
@@ II.1.1 Objective function
In the context of VQA, the specific choice of the Hamiltonian is not particularly significant. One can select any Hermitian operator and designate it as the Hamiltonian, then attempt to find the ground state. However, an essential property that the operator must possess is the ability to decompose into several operators that can be measured by the quantum processor. We will explore this concept in more detail in section II.1.3.
@@
$\langle H\rangle_{U(\theta)}\equiv\bra{0}U^{\dagger}(\theta)HU(\theta)\ket{0}~{}.$ (1)
It’s important to note that we initialize our qubits in the state |$\ket{0}$ and subsequently evolve them using the parameterized quantum circuit (PQC). As shown in Eqn. 1, the objective function depends on the unitary time evolution operator, making it a function of classical tunable parameters denoted as $\theta$. In some situations, it can be advantageous to utilize objective functions other than the Hamiltonian expectation value. Consequently, we have some flexibility in the specific form of the objective function, but we require that the objective function $\mathcal{O}$ be a function of $\theta$ and $\langle H\rangle_{U(\theta)}$, expressed as $\mathcal{O}\big{(}\theta~{},\{\langle H\rangle_{U(\theta)}\}\big{)}$.
@@
    The most natural approach to representing our Hamiltonian involves expressing it as a linear combination of basic tensor products of Pauli matrices: $\sigma_{i}$, with $\sigma_{i}$ taking values from the set $\{I,\sigma_{x},\sigma_{y},\sigma_{z}\}$. These tensor products are commonly referred to as Pauli strings, defined as $\hat{P}=\otimes_{j=1}^{n}\sigma_{i}^{(j)}$, where the index $j$ corresponds to different qubits. This representation is well-suited because the Pauli matrices $\sigma_{i}$ collectively form a basis for all Hermitian 2 by 2 matrices and serve as the generators of the special unitary group SU(2). The Hamiltonian, then, can be decomposed as $H=\sum_{k=1}^{M}c_{k}\hat{P}_{k}~{}.$ In this decomposition, the complex coefficient $c_{k}$ pertains to the k-th Pauli string, and the total number of Pauli strings, denoted as $M$, is contingent upon the specific form of the Hamiltonian under consideration.
@@
    Rather than optimizing with respect to the expectation value of an operator, certain VQAs necessitate a subroutine for optimizing the state obtained from the PQC $U(\theta)$, denoted as |$\ket{\Psi}_{U(\theta)}$, with respect to a target state, denoted as |$\ket{\Psi}$. We then define fidelity as the following: $F\big{(}\Psi,\Psi_{U(\theta)}\big{)}\equiv|\braket{\Psi}{\Psi_{U(\theta)}}|^{2}=\braket{\operatorname{Proj}_{\Psi}}_{U(\theta)}~{},$ which is the expectation value of the projection operator on the target state $\Psi$: $\operatorname{Proj_{\Psi}}=|\Psi\rangle\langle\Psi|~{}.$ Conventionally, VQA involves minimizing the objective function. In this context, it is natural to use infidelity, represented as $1-F\big{(}\Psi,\Psi_{U(\theta)}\big{)}$, or alternatively, the negative fidelity, denoted as $-\braket{\operatorname{Proj}_{\Psi}}_{U(\theta)}$.
    The expression of objective functions based on fidelities is commonly utilized in state preparation algorithms in quantum optics Kottmann et al. (2021); Krenn et al. (2020) and quantum machine learning Benedetti et al. (2019); Cheng et al. (2018); Huang et al. (2021b); Pérez-Salinas et al. (2020).
Additional objective functions can also be developed, and any cost function expressed operationally can be considered a viable option Bharti et al. (2021). Two examples include the Conditional-Value-at-Risk (CVaR) Barkoutsos et al. (2020) and the Gibbs objective function Li et al. (2020). These two forms can be simplified to the mean energy, denoted as $\braket{H}~{}.$ Importantly, their performance is at least as effective as utilizing the mean energy Li et al. (2020); Barkoutsos et al. (2020).
II.1.2 Parameterized quantum circuits (PQC)
Next, we delve into the second critical component of a VQA, which is the quantum circuit responsible for generating the state that optimally satisfies the defined objective. As described above, it prepares the state by an act of unitary that depends on classical parameters: $\theta~{},$ hence, it is called the parameterized quantum circuit: PQC.
The state prepared is defined as the following: |$\ket{\Psi(\theta)}=U(\theta)\ket{\Psi_{0}}~{},$ where |$\ket{\Psi_{0}}$ is some initial state, for example, |$\ket{0}^{\otimes n}~{},$ where $n$ is the number of qubits. As remarked earlier, $n$ should be quite small, due to the limited ability of the NISQ devices, we can only accommodate very shallow circuits. In some VQAs, it is convenient to use other initialization states. We can use a unitary operator (for example, a rotation operator) $P(\phi)$, and the initial state: |$\ket{\Psi_{0}}=P(\phi)\ket{0}^{\otimes n}~{}.$

### v2

# A comprehensive review of Quantum Machine Learning: from NISQ to Fault Tolerance

@@ 1 Introduction
Quantum machine learning represents a highly promising realm in contemporary physics and computer science research, with far-reaching implications spanning quantum chemistry [108], artificial intelligence [89], and even high-energy physics [7]. Nevertheless, it remains in its nascent stages of development. This is evident from the absence of a precise definition for quantum machine learning. Some describe it as the convergence of quantum computing and machine learning, wherein machine learning algorithms are executed on quantum devices. In simpler terms, it can be thought of as the quantum counterpart to classical machine learning.
@@
In recent times, artificial intelligence, exemplified by technologies like ChatGPT, has become an integral part of everyday life. It’s entirely plausible that, in the future, we will harness artificial intellegence in an even wider array of applications, including medical diagnostics, education, and aiding scientific research. Much of artificial intellegence’s success hinges on machine learning, a departure from traditional computer programs that entail crafting explicit instructions to directly solve problems [112]. Machine learning models, on the other hand, are trained on real-world data stored in a dataset (often denoted as $\mathcal{D}$), acquiring the ability to tackle problems autonomously. This represents a departure, in a sense, from the conventional von Neumann model of digital computing.
@@
Researchers might initially wonder why there’s a need for quantum machine learning when its classical counterpart has demonstrated impressive performance. One of the primary rationales stems from the fact that classical machine learning relies heavily on linear algebra procedures [18]. For instance, classical machine learning can address problems like identifying the optimal hyperplane to separate two data clusters by employing matrix inversion techniques. Quantum mechanics, at its core, is inherently grounded in linear algebra, where we construct a Hilbert space ($\mathcal{H}$) of a system described by an Hermitian operator known as the Hamiltonian ($\hat{H}$) [114]. Over time, it has become evident that quantum computing has the potential to dramatically enhance a computer’s problem-solving capabilities in certain specific scenarios. To illustrate, consider the task of matrix inversion again. Classical computers typically require computational efforts with a complexity of $\mathcal{O}(N\log{N})$ to accomplish this, while quantum computers can leverage algorithms like the Harrow-Hassidim-Lloyd algorithm (HHL) algorithm [58], which has a complexity of $\mathcal{O}((\log{N})^{2})$, leading to significant speedup, in certain conditions like sparsity, fault-tolerance, small condition numbers, and fast interfaces between classical and quantum processors. Although there is no known explicit realizations of HHL algorithm at the large scale at the moment, it is still an exciting direction deserves further studies, especially when quantum devices are fast developing with more and more capabilities.
@@
Due to its vague definition, quantum machine learning encompasses a broader spectrum of topics. For instance, quantum (shadow) tomography [2] has gained prominence, focusing on the characterization of a given quantum state by accumulating data from various measurements. This involves determining the minimum number of identical quantum state copies needed to extract sufficient information about the state’s properties. Another facet involves machine learning for quantum physics, which entails employing (perhaps classical) machine learning tools to explore various aspects of quantum physics. Additionally, some developments in both quantum algorithms and quantum hardware are often encompassed within the broader umbrella of quantum machine learning.
@@
The solution to this predicament is known as quantum error correction (QEC) code [113]. Think of QEC as a safeguard for quantum information. Typically, quantum information is lost once it’s measured, which becomes especially likely in noisy environments. However, information protected by QEC can persist if it remains undamaged within certain limits. It’s worth noting that all error correction codes have their constraints, implying that information will inevitably be lost if it’s severely damaged. Nevertheless, error correction provides a protective buffer zone against such losses.
@@
Our objective is to implement QEC codes across all quantum devices, ushering in an era of fault-tolerant quantum computing (FTQC) in the future. This trajectory parallels the history of classical computing, where, before the invention of classical error correction, scaling up and running useful algorithms on classical computers was also a formidable challenge [29]. Today, we can reliably operate classical computers everyday. Given the promising advancements in quantum error correction in recent times, our optimism about the future of quantum computing remains steadfast.
@@
However, when it comes to machine learning, classical machine learning doesn’t inherently reject noise [88]. The widely recognized learning algorithm, known as stochastic gradient descent, explicitly incorporates noise, and surprisingly, this noise addition actually enhances its performance. To grasp this concept, consider that noise can effectively steer us away from saddle points, offering an automatic mechanism for avoiding them. In a way, one can interpret this as machine learning’s ability to withstand and even benefit from noise. Consequently, this insight suggests that running certain machine learning algorithms on current (NISQ) quantum devices could have some significance. This prospect promises to enrich our present-day experiences significantly, especially considering the challenges associated with constructing QEC, which could take a few years to become fully integrated. Prior to entering the era of fault-tolerant quantum computing, we’ll have the opportunity to experiment with quantum devices and apply machine learning techniques, adding a vibrant dimension to our current scientific research.
@@
In addition to discussing the current applications of quantum capabilities for machine learning, we should also let our imaginations soar. The era of fault-tolerant quantum computing (FTQC) is a foreseeable future, and it’s crucial to look ahead at what challenges we can tackle with quantum machine learning on FTQC devices. One of the most renowned and valuable algorithms in this context is the Harrow-Hassidim-Lloyd (HHL) algorithm [58]. Additionally, there are other algorithms that can be deployed to address a range of problems, such as principal component analysis [18].
@@
Beside the optimistic future that quantum machine learning has, there are also a number of controversial issues with the subject. For example, some might argue that the variational quantum algorithm will not work in some circumstance. People working on quantum landscapes theory observed the famous barren plateau phenomena which leads to a kind of no-go theorem [27, 86]. It amounts to situations that are hard to find minimum when we optimizes our objective function by gradient descent with randomized variational circuits.
@@
Debates indeed exist regarding whether quantum algorithms can consistently deliver exponential speedups. Some argue that quantum speedup is only guaranteed when dealing with quantum information. When it comes to classical information, it’s conceivable to design classical machine learning models capable of achieving comparable average prediction accuracy [121, 66]. In such cases, the computational complexity difference between classical and quantum approaches may, at worst, be a modest polynomial factor. This debate underscores the importance of carefully assessing the specific problem and context when considering the potential advantages of quantum algorithms. However, the landscape shifts when the objective is to attain a low worst-case prediction error. In this scenario, it becomes feasible to achieve an exponential divergence in computational complexities between classical and quantum approaches. This underscores the potential superiority of quantum algorithms when stringent requirements for worst-case accuracy are in play. This phenomenon can be viewed as a manifestation of classical effects casting their “shadow” in the quantum realm. It should come as no surprise to individuals well-versed in quantum theory that quantum predictions tend to align with classical theory when dealing with systems possessing a high degree of freedom. This approach can also be applied to tackle a critical issue in variational quantum algorithms known as the “output problem,” often referred to as the classical shadow.
@@
Indeed, it is truly heartening and captivating to witness the evolution of quantum computing beyond celebrated algorithms like Shor’s factoring algorithm [117] and Grover search [50]. Algorithms like HHL [58] offer a versatile blueprint, shedding light on how quantum computers hold the potential to deliver significant acceleration for fundamental tasks like clustering, pattern-matching, and principal component analysis. However, as Scott Aaronson aptly noted in his paper [1], the pursuit of these speedups will still demand significant effort and ingenuity, as nature continues to present challenges that compel us to work diligently for these advancements.
@@
We have structured our review as follows. In section 2, we focus on the historical and ongoing developments in Quantum Machine Learning (QML) during the Noisy Intermediate-Scale Quantum (NISQ) era. Within this era, one of the central frameworks is the Variational Quantum Algorithm (VQA) [17, 44, 108], which we explore in detail in section 2.1. VQA comprises four key elements: the choice of an objective function (discussed in section 2.1.1), the employment of Parameterized Quantum Circuits (PQC) with adjustable parameters optimized by classical algorithms (covered in section 2.1.2), measurement strategies (explored in section 2.1.3), and the classical optimizer responsible for minimizing the objective function (explained in section 2.1.4). Subsequently, we provide insights into the construction of the Quantum Neural Tangent Kernel (QNTK) [87] in section 2.2, offering a theoretical foundation for quantum neural networks and an understanding of stochastic gradient descent dynamics from first principles. Finally, we address the issue of barren plateaus in section 2.3, approaching this topic through the lens of quantum landscape theory [27]. Additionally, we present an alternative formulation of laziness [86], providing another perspective on the same problem.
@@
In section 3, our focus shifts to quantum algorithms that have the potential for exponential speedup in the Fault-Tolerant Quantum Computing (FTQC) era. We begin by introducing Quantum Phase Estimation (QPE) [29] in section 3.1, where we also delve into the Quantum Principal Component Analysis (QPCA) program [91]. Subsequently, we present a counterpoint regarding the attainability of exponential speedup for programs such as the recommendation system in section 3.2. Moving on to section 3.3, we introduce the pivotal Harrow-Hassidim-Lloyd (HHL) algorithm [58], which opens up possibilities for various new quantum machine learning algorithms. In section 3.3.1, we illustrate a visionary perspective on the future of machine learning algorithms with the aid of Carleman linearization [90] and the HHL algorithm. Finally, in section 3.4, we introduce Quantum Random Access Memory (QRAM), focusing on theoretical designs [56], practical implementation efforts [54], and its necessity in certain algorithms [18].
@@
In section 4, we delve into various topics that amalgamate quantum principles with statistical learning theory. Our primary focus is on shadow tomography [2] in section 4.1, where we explore the motivation behind shadow tomography and delve into the construction of the theorem. A pivotal subject inspired by shadow tomography, the classical shadow formalism [41], is thoroughly discussed in section 4.2. Moving on to section 4.2.1, we examine the application of classical shadow as an efficient quantum-to-classical information converter [66]. In section 4.3, we then shift our attention to the applications of Quantum Machine Learning (QML) in the study of quantum data and quantum simulators [18].
@@
2 Noisy intermediate-scale quantum (NISQ) era
The concept of the NISQ era, which stands for Noisy Intermediate-Scale Quantum, was introduced by John Preskill in 2018 [110]. It’s essential to understand that NISQ is primarily a hardware-oriented definition and doesn’t inherently imply a specific time frame. In NISQ devices, quantum circuits are implemented, where all gates adhere to a predefined graph $G$, with qubits represented by nodes in this graph. These gates typically operate on one or two qubits. Due to the noise associated with each gate operation, NISQ algorithms are naturally constrained to shallow circuit depths [12].
On the other hand, “near-term” algorithms are those designed for quantum devices expected to be available in the next few years, and this term doesn’t explicitly refer to the absence of Quantum Error Correction (QEC). The notion of what constitutes the “near-term” can be subjective because different researchers may have varying opinions on how many years can be considered “near-term.” Predicting experimental progress in quantum computing is always challenging, and such predictions can be influenced by human biases. Algorithms developed for near-term hardware may become unfeasible if hardware advancements do not align with the algorithm’s experimental requirements [17].
The majority of existing NISQ algorithms leverage quantum computers’ capabilities through a hybrid quantum-classical approach. These algorithms involve offloading the classically challenging aspects of certain computations to a quantum computer while performing the remaining tasks on a suitably powerful classical device. They operate by iteratively adjusting the parameters of a parametrized quantum circuit, making them known as Variational Quantum Algorithms (VQA) ([42, 26]).
2.1 Variational quantum algorithm (VQA)
The concept of Variational Quantum Algorithms (VQA) was initially introduced to address certain quantum chemistry challenges, notably through the well-known Variational Quantum Eigensolver (VQE) [97, 108, 131]. Additionally, VQA includes the Quantum Approximate Optimization Algorithm (QAOA) [44], which is designed for solving combinatorial optimization problems. Interestingly, reference [5] shows that VQE combined with filtering operators can be more efficient and accurate than standard VQE and QAOA in the context of combinatorial optimization.
@@ 2.1.1 Objective function
In the context of VQA, the specific choice of the Hamiltonian is not particularly significant. One can select any Hermitian operator and designate it as the Hamiltonian, then attempt to find the ground state. However, an essential property that the operator must possess is the ability to decompose into several operators that can be measured by the quantum processor. We will explore this concept in more detail in section 2.1.3.
@@
$\langle H\rangle_{U(\theta)}\equiv\bra{0}U^{\dagger}(\theta)HU(\theta)\ket{0}\leavevmode\nobreak\ .$ (1)
It’s important to note that we initialize our qubits in the state |$\ket{0}$ and subsequently evolve them using the parameterized quantum circuit (PQC). As shown in Eqn. 1, the objective function depends on the unitary time evolution operator, making it a function of classical tunable parameters denoted as $\theta$. In some situations, it can be advantageous to utilize objective functions other than the Hamiltonian expectation value. Consequently, we have some flexibility in the specific form of the objective function, but we require that the objective function $\mathcal{O}$ be a function of $\theta$ and $\langle H\rangle_{U(\theta)}$, expressed as $\mathcal{O}\big{(}\theta\leavevmode\nobreak\ ,\{\langle H\rangle_{U(\theta)}\}\big{)}$.
@@
    The most natural approach to representing our Hamiltonian involves expressing it as a linear combination of basic tensor products of Pauli matrices: $\sigma_{i}$, with $\sigma_{i}$ taking values from the set $\{I,\sigma_{x},\sigma_{y},\sigma_{z}\}$. These tensor products are commonly referred to as Pauli strings, defined as $\hat{P}=\otimes_{j=1}^{n}\sigma_{i}^{(j)}$, where the index $j$ corresponds to different qubits. This representation is well-suited because the Pauli matrices $\sigma_{i}$ collectively form a basis for all Hermitian 2 by 2 matrices and serve as the generators of the special unitary group SU(2). The Hamiltonian, then, can be decomposed as $H=\sum_{k=1}^{M}c_{k}\hat{P}_{k}\leavevmode\nobreak\ .$ In this decomposition, the complex coefficient $c_{k}$ pertains to the k-th Pauli string, and the total number of Pauli strings, denoted as $M$, is contingent upon the specific form of the Hamiltonian under consideration.
@@
    Rather than optimizing with respect to the expectation value of an operator, certain VQAs necessitate a subroutine for optimizing the state obtained from the PQC $U(\theta)$, denoted as |$\ket{\Psi}_{U(\theta)}$, with respect to a target state, denoted as |$\ket{\Psi}$. We then define fidelity as the following: $F\big{(}\Psi,\Psi_{U(\theta)}\big{)}\equiv|\braket{\Psi}{\Psi_{U(\theta)}}|^{2}=\braket{\operatorname{Proj}_{\Psi}}_{U(\theta)}\leavevmode\nobreak\ ,$ which is the expectation value of the projection operator on the target state $\Psi$: $\operatorname{Proj_{\Psi}}=|\Psi\rangle\langle\Psi|\leavevmode\nobreak\ .$ Conventionally, VQA involves minimizing the objective function. In this context, it is natural to use infidelity, represented as $1-F\big{(}\Psi,\Psi_{U(\theta)}\big{)}$, or alternatively, the negative fidelity, denoted as $-\braket{\operatorname{Proj}_{\Psi}}_{U(\theta)}$.
    The expression of objective functions based on fidelities is commonly utilized in state preparation algorithms in quantum optics [76, 78] and quantum machine learning [15, 31, 64, 107].
Additional objective functions can also be developed, and any cost function expressed operationally can be considered a viable option [17]. Two examples include the Conditional-Value-at-Risk (CVaR) [13] and the Gibbs objective function [83]. These two forms can be simplified to the mean energy, denoted as $\braket{H}\leavevmode\nobreak\ .$ Importantly, their performance is at least as effective as utilizing the mean energy [83, 13]. The authors of [94] extend VQE to be applicable to a large class of nonlinear problems, including nonlinear differential equations, and show, for example, that quantum circuits can be exponentially more efficient (in terms of expressivity) than classical matrix product states.
2.1.2 Parameterized quantum circuits (PQC)
Next, we delve into the second critical component of a VQA, which is the quantum circuit responsible for generating the state that optimally satisfies the defined objective. As described above, it prepares the state by an act of unitary that depends on classical parameters: $\theta\leavevmode\nobreak\ ,$ hence, it is called the parameterized quantum circuit: PQC.
The state prepared is defined as the following: |$\ket{\Psi(\theta)}=U(\theta)\ket{\Psi_{0}}\leavevmode\nobreak\ ,$ where |$\ket{\Psi_{0}}$ is some initial state, for example, |$\ket{0}^{\otimes n}\leavevmode\nobreak\ ,$ where $n$ is the number of qubits. As remarked earlier, $n$ should be quite small, due to the limited ability of the NISQ devices, we can only accommodate very shallow circuits. In some VQAs, it is convenient to use other initialization states. We can use a unitary operator (for example, a rotation operator) $P(\phi)$, and the initial state: |$\ket{\Psi_{0}}=P(\phi)\ket{0}^{\otimes n}\leavevmode\nobreak\ .$